OpenStudy (anonymous):

The carbon monoxide level in a city is predicted to be 0.02x3/2 + 1 ppm (parts per million), where x is the population in thousands. In t years the population of the city is predicted to be x(t) = 12 + 2t thousand people. Therefore, in t years the carbon monoxide level will be P(t) = 0.02(12 + 2t)3/2 + 1 ppm. Find P'(12), the rate at which carbon monoxide pollution will be increasing in 12 years. P'(12) = I'm confused the derivative is .06(12+2(2t)^1/2+1 i thought and i got 1.36 when i plug in why is that wrong

OpenStudy (anonymous):

did you use the chain rule for that because its suppose to be (3/2)(.02)=.03 but then when you use chain rule you multiple 2(.03)=.06 because it said your answer was wrong too

OpenStudy (anwaarullah):

.06(12+2t)^.5 solve

OpenStudy (anwaarullah):

dnt add 1

OpenStudy (anonymous):

here this is my work and you can see if i went wrong somewhere but my teacher told us to use the chain rule for this problem so i know some certain parts are right u=.02u^3/2 = 3/2(.02)u^3/2-2/2=.03(u)^1/2 y=(2t+12)=2 (.03)=.06 =.06(12+2(12))^1/2 +1 ppm =.06(36)^1/2+1=.06(6)+1 =1.36

OpenStudy (anonymous):

why would you not add one??

OpenStudy (anwaarullah):

1 is constant ,,derivative of constan=0

OpenStudy (anonymous):

well .06(12+2(12)^.5 is wrong too

OpenStudy (anonymous):

and i can only submit one more time because so far the equations i been given are wrong is it possible that i need to convert my answer 1.36 to something because i been told that my work was actually right ??? thats why im really confused

OpenStudy (anwaarullah):

whats the answer btw

OpenStudy (anonymous):

but i did get told that the add one was suppose to be zero too but so that would make my answer equal with the work i did take off the add one =.36??

OpenStudy (anonymous):

nevermind my work was right i just needed to convert it and take off the one thank u

OpenStudy (anwaarullah):

yup...no problem,,,answer is .36 ,,,isnt

OpenStudy (anonymous):

i guess i needed to convert it to millions its webassign u know how funny online stuff can be

OpenStudy (anwaarullah):

yup,,i can understand lol